OpenStudy (anonymous):

find f(x) given f''(x) = 8 sqrtx

OpenStudy (anonymous):

you need more information than that!

OpenStudy (anonymous):

I know that I have to use antiderivatives and integration, but I don't know where to start.

OpenStudy (anonymous):

you have the second derivative only you can find the first derivative no problem use the power rule backwards

OpenStudy (kl0723):

can you integrate the first derivative to find f(x)?

OpenStudy (anonymous):

sure, but your answer will have a \(Cx\) in it for some unknown \(C\)

OpenStudy (anonymous):

f'(x) is sqrtx ^ 8 + C? I think that's the first derivative since if I work that in to the second it should give the second derivative.

OpenStudy (anonymous):

or am i way off...

OpenStudy (anonymous):

a bit off the anti derivative of \(x^{\frac{1}{2}}\) is \[\frac{x^{\frac{1}{2}+1}}{\frac{1}{2}+1}\] by using the powre ule backwards

OpenStudy (anonymous):

aka \(\large \frac{2}{3}x^{\frac{3}{2}}\)

OpenStudy (anonymous):

oh and i forgot your 8, so it should be \[f'(x)=\frac{16}{3}x^{\frac{3}{2}}+C\]

OpenStudy (aum):

\[\large f''(x) = 8\sqrt{x} = 8x^{1/2} \\ \large f'(x) = 8x^{3/2}*2/3 = 16/3 * x^{3/2} + C_1 \\ \large f(x) = 16/3 * x^{5/2} * 2/5 + C_1x + C_2 \\ \large f(x) = \frac{32}{15}x^{\frac 52} + C_1x + C_2 \]

OpenStudy (anonymous):

Oh man, ok thanks for the help you two. I will need to let this soak in and see if I can work this out as well on my own. Cheers.